# UAV-Assisted Real-Time Disaster Detection Using Optimized Transformer Model

### v1

@@ IV-E Visual Results Analysis
Figure 4 shows the visual results of the performance of our quantized model on our datasets. Usually, the model mistakes similar classes, such as landslide and mudslide, as seen from the last example in the third row. Moreover, the model produces mistakes due to its tendency to associate specific objects or features with particular classes. For example, the last picture in the first row shows scenery that can be associated with collapsed buildings. Finally, we observe failure cases when more than one disaster is illustrated in the photo. The last picture in the second row shows a bus fire, but cars are also present.
@@
Increasing the number of frames for each category could potentially solve the problems mentioned above. Moreover, introducing multilabel classification per frame might reduce some of the issues.
TABLE I: Summary of results of previous models trained on our datasets and our compressed, proposed model.
AIDER EmergencyNet [9] ARM Cortex-A57 0.957 25 0.09 0.78
TinyEmergencyNet [2] NVIDIA Quadro RTX 5000 GPU 0.940 - 0.04 0.15
MFEMANet [1] - 0.970 - - -
MobileNet [8] Jetson Nano 0.980 4 3.2 37.00
MobileNet Compressed [8] Jetson Nano - 71 3.2 7.00
Our Model FP16 NVIDIA T4 GPU 16 GB GPU 0.980 336.92 27.5 58.89
Our Model INT8 NVIDIA T4 GPU 16 GB GPU 0.977 236.41 27.5 36.76
Our Model INT8 Jetson Nano 0.977 45.22 27.5 34.00
DFAN DFAN [36] NVIDIA GPU 2070 12 GB GPU 0.870 70.55 23.9 83.63
DFAN Compressed [36] NVIDIA GPU 2070 12 GB GPU 0.860 125.33 23.9 41.09
MobileNet [36] NVIDIA GPU 2070 12 GB GPU 0.810 - 3.2 37.00
MAFire–Net [5] NVIDIA GeForce RTX-3090 GPU 0.875 78.31 - 74.43
ADFireNet [3] NVIDIA GeForce RTX-3090 GPU 0.900 72.50 7.2 38.00
Our Model FP16 NVIDIA T4 GPU 16 GB GPU 0.912 305.43 27.5 58.85
Our Model INT8 NVIDIA T4 GPU 16 GB GPU 0.910 240.17 27.5 34.91
Our Model INT8 Jetson Nano 0.910 45.04 27.5 34.00
@@ IV-F Ablation Experiments
In this section, we evaluate the performance of our model while employing different optimization techniques. Apart from the proposed TensorRT optimization, we use MinMax, EMA, OMSE, Percentile, and FQ-ViT quantization techniques. Results can be found in Table II. MinMax, EMA, OMSE, and Percentile optimization techniques were designed mainly for CNN architectures. Therefore, they focus on Convolutional, Linear, and MatMul modules while lacking solutions for handling Softmax, LayerNorm, and GELU layers. This means many layers stay in a 32-bit floating point, resulting in the mix-precision network and unchanged model size; however, a significant speedup, enough for real-time performance, can be seen for all these methods. FQ-ViT, specially tailored for transformer-based architectures, introduces powers-of-two scale quantization and log-int quantization for LayerNorm and Softmax, respectively. However, FQ-ViT does not focus on GELU layers, leading to partial performance. Due to dequantized floating-point parameters during inference, efficient low-precision arithmetic units are not fully utilized, and, therefore, we get unsatisfactory model acceleration.
TABLE II: Optimization methods used on the proposed model.
@@ IV-G Comparison With the State-of-the-Art
Table I shows the performances of state-of-the-art models and our proposed method. Our model outperforms almost all previously trained networks in terms of f1-score on the AIDER dataset. In [8], authors used class weights due to unbalanced data to train MobileNet, while our results purely rely on transformers’ capability to capture valuable information from frames. The limited network, such as MobileNet, performs poorly when trained on the DFAN dataset, which contains few samples and many classes. At the same time, our model achieves the highest performance in terms of f1-score, outperforms other solutions [36], [5], [3], [4] on high computational GPU devices regarding speed and achieves real-time performance on resource-constrained devices. Consequently, our proposed framework’s scalability and adaptability make it a valuable tool for deploying real-time disaster detection solutions on resource-limited UAV platforms.
@@ V Conclusion
This work confirms the effectiveness of transformer-based architectures for real-time disaster classification on UAVs. By utilizing advanced quantization techniques, the model substantially reduces memory footprint and latency while maintaining accuracy across various disaster types. The results show that INT8 quantization with TensorRT significantly enhances model performance on edge devices, with minimal impact on accuracy for distinctive classes. The proposed framework demonstrates robust classification capabilities in diverse disaster scenarios and highlights the potential of deploying transformer models on UAVs for prompt and autonomous disaster response. Future work will focus on extending the framework with additional disaster scenarios and expanding the DisasterEye dataset.

### v2

@@ IV-E Visual Results Analysis
Fig. 4 shows the visual results of the performance of our quantized model on our datasets.
@@
Usually, the model mistakes similar classes, such as landslide and mudslide, as seen from the last example in the third row. Moreover, the model makes errors because of its tendency to associate specific objects or features with particular classes. For example, the last picture in the first row shows scenery that can be associated with collapsed buildings. Finally, we observe failure cases when more than one disaster is illustrated in the photo. The last picture in the second row shows a bus fire, but cars are also present. Increasing the number of frames for each category could potentially solve the problems mentioned above. Moreover, introducing multilabel classification per frame might reduce some of the issues.
@@ IV-F Ablation Experiments
In this section, we evaluate the performance of our model while employing different optimization techniques. In addition to the proposed TensorRT optimization, we use MinMax, EMA, OMSE, Percentile, and FQ-ViT quantization techniques. The results can be found in Table I.
MinMax, EMA, OMSE, and Percentile optimization techniques were designed mainly for CNN architectures. Therefore, they focus on Convolutional, Linear, and MatMul modules while lacking solutions for handling Softmax, LayerNorm, and GELU layers. This means many layers stay in a 32-bit floating point, resulting in the mix-precision network and unchanged model size; however, a significant speed-up, enough for real-time performance, can be seen for all these methods.
FQ-ViT, specially tailored for transformer-based architectures, introduces powers-of-two scale quantization and log-int quantization for LayerNorm and Softmax, respectively. However, FQ-ViT does not focus on GELU layers, leading to partial performance. Due to dequantized floating-point parameters during inference, efficient low-precision arithmetic units are not fully utilized, and, therefore, we get unsatisfactory model acceleration.
TABLE I: Optimization methods used on the proposed model.
@@ IV-G Comparison With the State-of-the-Art
Table II shows the performance of the state-of-the-art models and our proposed method. Our model outperforms almost all previously trained networks in terms of f1-score on the AIDER dataset. In [8], the authors used class weights due to unbalanced data when training their MobileNet. This helped the model make better predictions. Our results, on the other hand, rely purely on the transformers’ capability to capture valuable information from frames. Regarding FPS, our proposed model surpasses all lighter models that contain fewer parameters and smaller model sizes. Finally, it demonstrates real-time performance on resource-constrained devices such as the Jetson Nano.
The limited network, such as MobileNet, performs poorly when trained on the DFAN dataset. At the same time, our model achieves the highest performance in terms of f1 score. It outperforms other solutions such as [38], [5], [3], [4] on high computational GPU devices regarding speed and achieves real-time performance on resource-constrained devices.
These results show that our proposed framework’s scalability and adaptability make it a valuable tool for deploying real-time disaster detection solutions on resource-limited UAV platforms.
TABLE II: Summary of results of previous models trained on our datasets and our compressed, proposed model.
AIDER EmergencyNet [9] ARM Cortex-A57 0.957 25 0.09 0.78
TinyEmergencyNet [2] NVIDIA Quadro RTX 5000 GPU 0.940 - 0.04 0.15
MFEMANet [1] - 0.970 - - -
MobileNet [8] Jetson Nano 0.980 4 3.2 37.00
MobileNet Compressed [8] Jetson Nano - 71 3.2 7.00
Our Model FP16 NVIDIA T4 GPU 16 GB GPU 0.980 336.92 27.5 58.89
Our Model INT8 NVIDIA T4 GPU 16 GB GPU 0.977 236.41 27.5 36.76
Our Model INT8 Jetson Nano 0.977 45.22 27.5 34.00
DFAN DFAN [38] NVIDIA GPU 2070 12 GB GPU 0.870 70.55 23.9 83.63
DFAN Compressed [38] NVIDIA GPU 2070 12 GB GPU 0.860 125.33 23.9 41.09
MobileNet [38] NVIDIA GPU 2070 12 GB GPU 0.810 - 3.2 37.00
MAFire–Net [5] NVIDIA GeForce RTX-3090 GPU 0.875 78.31 - 74.43
ADFireNet [3] NVIDIA GeForce RTX-3090 GPU 0.900 72.50 7.2 38.00
Our Model FP16 NVIDIA T4 GPU 16 GB GPU 0.912 305.43 27.5 58.85
Our Model INT8 NVIDIA T4 GPU 16 GB GPU 0.910 240.17 27.5 34.91
Our Model INT8 Jetson Nano 0.910 45.04 27.5 34.00
@@ V Conclusion
This work confirms the effectiveness of optimized transformer-based architectures for real-time disaster classification on UAVs. By utilizing advanced quantization techniques, the model substantially reduces memory footprint and latency while maintaining accuracy across various disaster types.
The results show that the precision calibration of FP16 and INT8 in TensorRT gives real-time performance with increased FPS compared to the original model and reduced model size by 50% and 70%, respectively. The optimized proposed model maintains high accuracy when datasets have sufficient samples per category, such as AIDER, while datasets with many classes but limited samples per category, such as DFAN, may experience some accuracy drop. Finally, TensorRT significantly enhances the model performance on edge devices such as Jetson Nano, achieving real-time performance.
The proposed framework demonstrates robust classification capabilities in diverse disaster scenarios and highlights the potential of deploying transformer models on UAVs for a prompt and autonomous disaster response. In addition, the performance of the proposed model on our DisasterEye dataset highlights its potential as a benchmark for the classification of real-life disaster scenarios. Future work will focus on extending the framework by adding more disaster scenarios and expanding the DisasterEye dataset.